OpenStudy (anonymous):

Braden's family had 1 and one-third gallons of milk in the refrigerator before dinner. After dinner, they had finished three-eighths of the milk. How many gallons of milk did they drink at dinner?

OpenStudy (anonymous):

1 1/3 gallons = 4/3 gallons 3/8 of 4/3 = 3/8 * 4/3 = 12/24 = 1/2 They drank 1/2 gallon

OpenStudy (anonymous):

no they didn't

OpenStudy (anonymous):

(3/8) / (1/2) = (5/8) / x cross multiply 3/8x = (1/2)(5/8) 3/8x = 5/16 x = (5/16) / (3/8) x = 5/16 * 8/3 x = 40/48 x = 5/6 so 3/8 is 1/2 5/8 = 5/6 1/2 + 5/6 = 4/3 3/6 + 5/6 = 8/6 8/6 = 8/6 (correct) They drank 1/2 gallons leaving 5/6 gallons left.

OpenStudy (anonymous):

@wio ...can you take a look at this. Am I doing something wrong

OpenStudy (anonymous):

Let's see... well

OpenStudy (anonymous):

Your answer looks correct.

OpenStudy (anonymous):

I would write it as: \(4/3 \cdot (1-3/8)\).

OpenStudy (anonymous):

but it would still be 1/2 gallons drank...right ?

OpenStudy (anonymous):

basically, it is just asking what is 3/8 of 1 1/3 gallons..right ?

OpenStudy (anonymous):

oh wait, never mind... it is just 1/2